# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ 5 Output
  • $|S|=n$ (number of signers)
$R_{2}^{\epsilon}$ $\leq 1+\frac{n^{2}\cdot\beta^{2}}{(2\gamma_{1}/n)^{2}}=1+\frac{n^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
$=1+\frac{n^{4}\cdot 196^{2}}{4\cdot(524288)^{2}}=1+\frac{n^{4}\cdot 38416}{1.099\times 10^{12}}$
Concrete values (per-coordinate): Since nonce coordinates are sampled independently, the full-vector Rényi divergence is $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ where $n\ell=1280$ for ML-DSA-65:
$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}-1$ $R_{2}^{\mathsf{vec}}=(R_{2}^{\mathsf{coord}})^{1280}$ $\sqrt{R_{2}^{\mathsf{vec}}}$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$
@@
The tail probability $\epsilon$ satisfies [Raccoon2024, Lemma 4.2]:
$\epsilon<\left(\frac{2e\cdot\beta}{N}\right)^{n}/\sqrt{2\pi n}$
For $|S|=17$: $\epsilon<(2e\cdot 196/61681)^{17}/\sqrt{34\pi}<10^{-30}$.
@@
Conclusion: Applying Lemma Q.12 with the full-vector Rényi divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$, and taking a union bound over $q_{S}$ queries for the smoothing cost:
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}}+q_{S}\cdot\epsilon$
Security interpretation: For ML-DSA-65 with $\mathsf{Adv}_{\mathsf{U}}\leq 2^{-192}$ (NIST Level 3), $\epsilon<10^{-30}\approx 2^{-100}$, and $q_{S}\leq 2^{30}$:
$|S|$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$ Total bound
$\leq 17$ $\approx 2^{-93.3}$ $\approx 2^{-70}$ ($q_{S}\cdot\epsilon$ dominant)
$\leq 25$ $\approx 2^{-83.5}$ $\approx 2^{-70}$ ($q_{S}\cdot\epsilon$ dominant)
$\leq 33$ $\approx 2^{-58.5}$ $\approx 2^{-58.5}$ ($\sqrt{R_{2}^{\mathsf{vec}}}$ dominant)
For $|S|\leq 25$, the $q_{S}\cdot\epsilon$ term dominates and the Rényi composition cost is modest (<13<13 bits). For $|S|\geq 25$, the full-vector $\sqrt{R_{2}^{\mathsf{vec}}}$ term becomes significant due to the $(1+\delta)^{1280}$ growth.

### v5

@@ 5 Output
  • $m=|S|$ (signing set size; note $n=256$ is the ring degree)
$R_{2}^{\epsilon}$ $\leq 1+\frac{m^{2}\cdot\beta^{2}}{(2\gamma_{1}/m)^{2}}=1+\frac{m^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
$=1+\frac{|S|^{4}\cdot 196^{2}}{4\cdot(524288)^{2}}=1+\frac{|S|^{4}\cdot 38416}{1.099\times 10^{12}}$
Why this bound is conservative ($\approx|S|^{3}/12$ looser than the Gaussian CLT): The Raccoon Lemma 4.2 formula $m^{2}\delta^{2}/B^{2}$ uses a moment bound that treats the IH sum as if its worst-case chi-squared scales with $m^{2}/B^{2}$ (reflecting a worst-case uniform-distribution approximation over the aggregate range $mB/2$). The Gaussian CLT-based Fisher information bound: the IH variance is $\mathrm{Var}(Y)=\gamma_{1}^{2}/(3m)$ (sum of $m$ i.i.d. uniforms on $\gamma_{1}/m,\gamma_{1}/m]$ with variance $(\gamma_{1}/m)^{2}/3$), giving Fisher information $I_{F}=3m/\gamma_{1}^{2}$ and $\chi^{2}\approx\delta^{2}\cdot I_{F}=3m\delta^{2}/\gamma_{1}^{2}$. Substituting $\gamma_{1}=Bm/2$ (from $B=2\gamma_{1}/m$): $\chi^{2}=3m\delta^{2}/(Bm/2)^{2}=12\delta^{2}/(mB^{2})$ (proportional to $m^{-1}$, since variance grows as $m$), while the Raccoon formula gives $m^{2}\delta^{2}/B^{2}$ (proportional to $m^{2}$). The ratio is
$\frac{m^{2}\delta^{2}/B^{2}}{12\delta^{2}/(mB^{2})}=\frac{m^{3}}{12}\approx\frac{|S|^{3}}{12},$
which equals $409$ for $|S|=17$ (the Gaussian approximation; the exact D2 ratio from FFT values is ${\approx}441$, cited as ${\approx}440\times$ throughout the paper). The Raccoon formula is formally correct as an upper bound; the direct Gaussian-based analysis (Lemma Q.22) gives the tight bound by using the actual IH variance rather than the aggregate-range worst case.
Concrete values (per-coordinate): Since nonce coordinates are sampled independently before rejection sampling, the full-vector Rényi divergence factorizes as $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ where $n\ell=1280$ for ML-DSA-65 (here $n=256$ is the ring degree, $\ell=5$ is the secret dimension). This factorization applies to the pre-rejection distributions $P$ and $Q$. To handle rejection sampling: the smooth Rényi framework removes the tail region $\{|Y|>\gamma_{1}-\beta\}$ from the support of $P$, producing a restricted distribution $P^{\prime}$ supported on the same acceptance set $\{|Y|\leq\gamma_{1}-\beta\}$ as $Q^{\prime}$. Within this support the rejection condition is deterministic (every sample is accepted), so the factorization carries through for the accepted distributions $P^{\prime}$ and $Q^{\prime}$. The $\epsilon$ term in the bound accounts for the tail mass removed from $P$; see [Raccoon2024] for the precise smooth Rényi treatment. The resulting bound is:
$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}-1$ $R_{2}^{\mathsf{vec}}$ $\sqrt{R_{2}^{\mathsf{vec}}}$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$
@@
The tail probability $\epsilon$ satisfies [Raccoon2024, Lemma 4.2] (using $m=|S|$ and $B=2\lfloor\gamma_{1}/|S|\rfloor$):
$\epsilon<\left(\frac{2e\cdot\beta}{B}\right)^{m}/\sqrt{2\pi m}$
For $|S|=17$: $B=2\lfloor 524288/17\rfloor=2\times 30840=61680$, so $\epsilon<(2e\cdot 196/61680)^{17}/\sqrt{34\pi}<10^{-30}$.
@@
Conclusion: Applying Lemma Q.13 to the joint distribution over $q_{s}$ independent signing sessions (joint Rényi divergence $(R_{2}^{\mathsf{vec}})^{q_{s}}$) and taking a union bound over the smoothing tails:
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}\leq\bigl(R_{2}^{\mathsf{vec}}\bigr)^{q_{s}/2}\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}}+q_{s}\cdot\epsilon$
Security interpretation (direct bound, Corollary Q.24): Using $R_{2}^{\mathsf{vec}}\approx(1+\chi^{2}_{\mathsf{direct}})^{1280}$ where $\chi^{2}_{\mathsf{direct}}$ is the directly computed chi-squared divergence (Corollary Q.24; slightly smaller than the Gaussian approximation $3|S|\beta^{2}/\gamma_{1}^{2}$), the multi-session factor contributes $\approx\frac{q_{s}\cdot n\ell}{2}\cdot\log_{2}(1+\chi^{2}_{\mathsf{direct}})$ bits of security loss. For ML-DSA-65 with $\mathsf{Adv}_{\mathsf{U}}\leq 2^{-192}$:
$|S|$ $\chi^{2}_{\mathsf{direct}}$ Loss/query (bits) Max $q_{s}$ for $\leq$10-bit loss
44 $8.4\times 10^{-7}$ $7.7\times 10^{-4}$ $\approx 2^{14}$
$17$ $6.6\times 10^{-6}$ $6.1\times 10^{-3}$ $\approx 1600$
$33$ $1.4\times 10^{-5}$ $1.2\times 10^{-2}$ $\approx 830$
The conservative bound (Raccoon Lemma 4.2, $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ for $|S|=17$) is vacuous for $q_{s}\gtrsim 36$ queries; the direct bound above should be used for concrete security estimates.